OpenStudy (jessie125):

A cylinder has a radius of 10 cm and a height of 8 cm. What is the volume of the cylinder? A. 160 straight pi cm3 B. 480 straight pi cm3 C. 640 straight pi cm3 D. 800 straight pi cm3

satellite73 (satellite73):

i think it is \[\pi \times r^2\times h\] for the volume area of the circle times the height

satellite73 (satellite73):

so \[\pi\times 10^2\times 8\] in your case

OpenStudy (jessie125):

so c

satellite73 (satellite73):

no

satellite73 (satellite73):

square the ten, not the eight

OpenStudy (jessie125):

oh so B

satellite73 (satellite73):

what is \(10^2\)??

OpenStudy (jessie125):

10x10

satellite73 (satellite73):

=?

OpenStudy (jessie125):

100=800 thx

satellite73 (satellite73):

yw btw that has to be the cutest picture i have ever seen

OpenStudy (jessie125):

thank you